# Labrador: Exploring the limits of masked language modeling for laboratory data

### v1

@@ 6 Conclusion
In this work, we presented Labrador, a novel Transformer architecture that can be pre-trained on (integer, float) data such that the float values are treated in a continuous manner. We showed that Labrador outperforms BERT across all downstream evaluations and appears to learn monotonic axes corresponding to lab value in its embedding space. However, despite the success seen in pre-training, both models fail to outperform XGBoost on four outcome prediction tasks. Our ablation studies reveal that transfer learning largely fails in this setting, despite being strictly more successful for Labrador than for BERT. We hypothesize that this failure is largely due to a lack of pre-training data scale and an insufficient characterization of the data generating process underlying each patient using labs in isolation from other clinical data modalities. As a result, we urge future work to focus on harmonizing disparate datasets and jointly modeling several categories of EHR data.
A persistent challenge in medical machine learning for structured data is the lack of universally accepted evaluations. Structured data predictions are fundamentally difficult to verify, as there is no available human baseline for performance. Although physicians take structured data into account when assessing patients, they do not perform regression or classification using these data in isolation. As a result, we have a poor sense of the Bayes error rate for these types of prediction tasks. We argue that a prerequisite to progressing the self-supervised learning efforts on EHR data is the establishment of universally accepted evaluations that will provide a clear picture of the field’s progress over time in comparison to tree-based methods (Bellamy et al., 2020).
7 Code and data availability
The MIMIC-IV database can be accessed via PhysioNet after completing a short online training. The lab data that was used to pre-train Labrador and BERT come from the labevents table within MIMIC-IV and our pre-processing code can be found in our codebase at https://github.com/DavidBellamy/labrador. The COVID-19, cancer diagnosis and alcohol consumption fine-tuning datasets are public and therefore we provide our pre-processed versions of these datasets directly in our GitHub repository under the data/ directory.
The GitHub repository also contains all of the code necessary to replicate the pre-training data pre-processing, model pre-training, fine-tuning and other evaluations. We cannot directly share the sepsis dataset, however with access to MIMIC-IV the steps described in Section C.3.2 recreate our evaluation dataset. Finally, the weights for the pre-trained Labrador and BERT models can be downloaded from HuggingFace at https://huggingface.co/Drbellamy/labrador.
8 Acknowledgements
DB would like to thank Dr. Tianxi Cai for her insightful idea to transform the lab values using their empirical CDF. We also thank Dr. Leo Celi for his help teaching us about the invaluable MIMIC dataset and the importance of advancing this work.
9 Author contributions
BK helped write the UMAP code and the manuscript, CW helped write the fine-tuning experiment code, DB wrote the remaining code and conducted the experiments present in the paper, and AB helped oversee the project and conceive of the original idea. Correspondence to David Bellamy.
@@ A.2 Masked language models
Appendix B Parameter-matched BERT experiments
@@
Appendix C Further experimental details
C.1 UMAP
After performing several scaling experiments with UMAP reduction on the full test split of the pre-training data, we observed that the results were not significantly altered if we used a random sub-sample of the data. We randomly selected bags of labs totalling 50,000 (lab code, lab value) pairs from the test split. These bags separately underwent Labrador and BERT’s forward pass to obtain the final layer embeddings from each Transformer. The UMAP algorithm was run with n_neighbors=600, min_dist=0.99 and a random seed of 3141592. In preliminary experiments, we found that increasing the number of neighbors and increasing the minimum distance between points in the reduction provided a better global structure of the embedding space. In consultation with the MIMIC database authors, we manually labeled lab codes according to their panel membership, such as the Complete Blood Count (see Appendix F for all panels). This allowed us to color Figure 2 accordingly.
C.2 Intrinsic imputations
@@
C.3 Pre-processing the evaluation data
C.3.1 COVID-19 cohort
For COVID-19 diagnosis, we used a public dataset (Cabitza et al., 2020; 2021) that contains blood work for 1,624 patients (52% COVID-19 positive) admitted at San Raphael Hospital from February to May 2020. The data contain 34 features and a binary target variable for COVID-19 diagnosis determined by RT-PCR at a later point in time. Features ‘CK’ (creatine kinase) and ‘UREA’ (blood urea nitrogen) were dropped from the data because their rates of missingness were too high (60% and 38.5%, respectively). Any patients that were missing greater than 25% of the remaining 32 features were dropped as well since we believe that the imputation quality for these patients would be too poor. This resulted in 1,312 patients.
@@ C.3.1 COVID-19 cohort
C.3.2 Sepsis cohort
We downloaded the MIMIC-IV sepsis cohort from Google BigQuery under physionet-data/mimic_derived/sepsis3. Accessing the MIMIC-IV data (Johnson et al., 2020; 2023) requires signing a data use agreement and completing a short online training, so we cannot share this dataset directly. However, after obtaining MIMIC-IV access on PhysioNet, you can find instructions on accessing the data using Google BigQuery.
@@ C.3.2 Sepsis cohort
C.3.3 Cancer diagnosis cohort
For cancer diagnosis, we used a public dataset shared by Tsai et al. (2022) that contains blood work for 1,336 patients. Briefly, each patient received a diagnosis of either cystitis, bladder cancer, kidney cancer, uterine cancer, or prostate cancer. The original dataset contained 39 features in addition to the 5-class target variable. Five of these 39 features were removed for having too much missingness (urine epithelium, A/G ratio, urine ketone, urine glucose, and strip WBC). Among the remaining 34 features, 24 were lab tests within the MIMIC-IV vocabulary, 2 lab tests were out of vocabulary, and there were 8 social variables (age, gender, smoking status, diabetes status, etc.). We manually verified that each in-vocabulary lab test was using the same units of measurement as in MIMIC-IV. For categorical variables that had missing values, we used the missing indicator method. We transformed all lab values using the eCDF method and missing lab values were handled in the same manner as for the COVID-19 dataset (Section C.3.1).
C.3.4 Alcohol consumption cohort
@@
C.4 Transformer fine-tuning
@@
def LabradorFinetuneWrapper(inputs: Dict[Tensor]) -> Tensor:
    x = base_model(inputs[‘lab_features’])
@@
    if ‘non_mimic_features’ in inputs.keys():
        non_mimic_features = dense(inputs[‘non_mimic_features’])
        x = concat([x, non_mimic_features], axis=1)
@@
C.5 Baseline methods for outcome prediction
@@
Appendix D Pre-training loss curves
@@
Appendix E Lab value correlation plot
@@
Figure 8: Lab value correlation matrix. A 70 x 70 matrix of Pearson correlation coefficients between each pair of the 70 most frequently ordered lab tests in MIMIC-IV. The 70 tests are grouped by the panel that they are most frequently ordered with (see Appendix F for panel definitions). Black lines divide the matrix into panel-by-panel correlation blocks.
Appendix F Lab panel definitions

### v2

@@ 6 Conclusion
In this work, we presented Labrador, a novel Transformer architecture that can be pre-trained on (integer, float) data such that the float values are treated in a continuous manner. We showed that Labrador outperforms BERT across all downstream evaluations and appears to learn monotonic axes corresponding to lab value in its embedding space. However, despite the success seen in pre-training, both models fail to outperform XGBoost on four outcome prediction tasks. Our ablation studies reveal that transfer learning largely fails in this setting despite being more successful for Labrador than for BERT. We hypothesize that this failure is largely due to a lack of pre-training data scale and an insufficient characterization of the data generating process underlying each patient using labs in isolation from other clinical data types. As a result, we urge future work to focus on harmonizing disparate datasets and jointly modeling several categories of EHR data.
A persistent challenge in medical machine learning for structured data is the lack of universally accepted evaluations. Model predictions are difficult to verify, since there is no human baseline for performance. As a result, we have a poor sense of the Bayes error rate for these types of prediction tasks. We argue that a prerequisite to progressing the self-supervised learning efforts on EHR data is the establishment of universally accepted evaluations that will provide a clear picture of the field’s progress over time in comparison to tree-based methods (Bellamy et al., 2020).
@@ A.2 Masked language models
Appendix B Pseudo-code Implementations
The continuous embedding layer is implemented as follows:
def ContinuousEmbeddingLayer(
    lab_values, token_embeddings):
    x = Dense_pw(lab_values,
        out_size=d_model,
        activation=‘linear’)
    x = x + token_embeddings
    x = Dense_pw(x, out_size=d_model,
        activation=‘relu’)
    x = LayerNorm(x)
    return x
The continuous prediction head is implemented as follows:
def ContinuousPredictionHead(
    final_layer_embeddings, probs):
    x = concat([final_layer_embeddings,
                probs])
    x = Dense_pw(x,
        out_size=embed_dim + probs_dim,
        activation=‘relu’)
    x = Dense_pw(x, out_size=1,
        activation=‘sigmoid’)
    return x
Where probs is the categorical head’s probability distribution over the lab code vocabulary. This supplies the continuous prediction head with each token’s predicted probability when attempting to impute the masked lab value.
Appendix C Embedding Space Visualization
Figure 2: UMAP Visualization of Labrador and BERT Embeddings. A. A global view of the embedding space structure for Labrador (left) and BERT (right). The 70 most frequently ordered lab tests are shown and colored according to the panel of tests they are typically ordered with (see Appendix I for all panel definitions). All labs on this figure are from the test split and were not seen during pre-training. B. Visualization of embeddings for four routinely collected laboratory measurements colored by lab value and scaled to the interval [0, 1]. Labrador appears to encode the measured lab value in a much more natural way with a smooth gradient for lab value compared to BERT.
Appendix D Imputation Evaluation of Labrador and BERT
Appendix E Parameter-matched BERT experiments
@@
Appendix F Further experimental details
F.1 UMAP
After performing several scaling experiments with UMAP reduction on the full test split of the pre-training data, we observed that the results were not significantly altered if we used a random sub-sample of the data. We randomly selected bags of labs totalling 50,000 (lab code, lab value) pairs from the test split. These bags separately underwent Labrador and BERT’s forward pass to obtain the final layer embeddings from each Transformer. The UMAP algorithm was run with n_neighbors=600, min_dist=0.99 and a random seed of 3141592. In preliminary experiments, we found that increasing the number of neighbors and increasing the minimum distance between points in the reduction provided a better global structure of the embedding space. In consultation with the MIMIC database authors, we manually labeled lab codes according to their panel membership, such as the Complete Blood Count (see Appendix I for all panels). This allowed us to color Figure 2 accordingly.
F.2 Intrinsic imputations
@@
F.3 Pre-processing the evaluation data
F.3.1 COVID-19 cohort
For COVID-19 diagnosis, we used a public dataset (Cabitza et al., 2020, 2021) that contains blood work for 1,624 patients (52% COVID-19 positive) admitted at San Raphael Hospital from February to May 2020. The data contain 34 features and a binary target variable for COVID-19 diagnosis determined by RT-PCR at a later point in time. Features ‘CK’ (creatine kinase) and ‘UREA’ (blood urea nitrogen) were dropped from the data because their rates of missingness were too high (60% and 38.5%, respectively). Any patients that were missing greater than 25% of the remaining 32 features were dropped as well since we believe that the imputation quality for these patients would be too poor. This resulted in 1,312 patients.
@@ F.3.1 COVID-19 cohort
F.3.2 Sepsis cohort
We downloaded the MIMIC-IV sepsis cohort from Google BigQuery under physionet-data/mimic_derived/sepsis3. Accessing the MIMIC-IV data (Johnson et al., 2020, 2023) requires signing a data use agreement and completing a short online training, so we cannot share this dataset directly. However, after obtaining MIMIC-IV access on PhysioNet, you can find instructions on accessing the data using Google BigQuery.
@@ F.3.2 Sepsis cohort
F.3.3 Cancer diagnosis cohort
For cancer diagnosis, we used a public dataset shared by Tsai et al. (2022) that contains blood work for 1,336 patients. Briefly, each patient received a diagnosis of either cystitis, bladder cancer, kidney cancer, uterine cancer, or prostate cancer. The original dataset contained 39 features in addition to the 5-class target variable. Five of these 39 features were removed for having too much missingness (urine epithelium, A/G ratio, urine ketone, urine glucose, and strip WBC). Among the remaining 34 features, 24 were lab tests within the MIMIC-IV vocabulary, 2 lab tests were out of vocabulary, and there were 8 social variables (age, gender, smoking status, diabetes status, etc.). We manually verified that each in-vocabulary lab test was using the same units of measurement as in MIMIC-IV. For categorical variables that had missing values, we used the missing indicator method. We transformed all lab values using the eCDF method and missing lab values were handled in the same manner as for the COVID-19 dataset (Section F.3.1).
F.3.4 Alcohol consumption cohort
@@
F.4 Transformer fine-tuning
@@
def LabradorFinetuneWrapper(
    inputs: Dict[Tensor]) -> Tensor:
    x = base_model(inputs[’lab_features’])
@@
    if ’non_mimic_features’ in inputs.keys():
        non_mimic_features = dense(
            inputs[’non_mimic_features’])
        x = concat([x, non_mimic_features],
                   axis=1)
@@
F.5 Baseline methods for outcome prediction
@@
Appendix G Pre-training loss curves
@@
Appendix H Lab value correlation plot
@@
Figure 8: Lab value correlation matrix. A 70 x 70 matrix of Pearson correlation coefficients between each pair of the 70 most frequently ordered lab tests in MIMIC-IV. The 70 tests are grouped by the panel that they are most frequently ordered with (see Appendix I for panel definitions). Black lines divide the matrix into panel-by-panel correlation blocks.
Appendix I Lab panel definitions